OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Plug the slope and point into point-slope form and simplify. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope.

OpenStudy (anonymous):

dude lol you should give igreen a medal!

OpenStudy (anonymous):

y=2x+4

OpenStudy (igreen):

Not quite.

OpenStudy (igreen):

\(y - 4 = 2(x - 3)\) Distribute 2 into the parenthesis: \(y - 4 = 2x - 6\) Add 4 to both sides, what's -6 + 4?

OpenStudy (anonymous):

-2